Use the information below for Nuthatch Corporation to answer the question that follow.

Nuthatch Corporation began its operations on September 1 of the current year. Budgeted sales for the first three months of business—September, October, and November—are $232,000, $307,000, and $424,000, respectively. The company expects to sell 30% of its merchandise for cash. Of sales on account, 80% are expected to be collected in the month of the sale and 20% in the month following the sale. The cash collections expected in November from accounts receivable are projected to be
Answer #1

correct option is "b"-280420

If 30% of merchandise sold is for cash then remaining 100-30 =70% if on credit .

Cash collections in November
Credit sales collections
October sales (307000*70%= 214900*20%) 42980
November sales (424000*70%= 296800*80%) 237440
Total cash collections in November from account receivable 280420